# Scaling Relations in the Phase-Space Structure of Dark Matter Haloes

### v1

@@ 6 Summary and Conclusions
Based on the comparisons in Figs. 24 and Table 1, we suggest the following two efficient methods to approximate the energy distributions and DFs of haloes. For simulated haloes, one can use the scaling relations in Eqs. (20) and (21) after obtaining the best-fit NFW profile and the numerical potential. In more approximate contexts, one can use the potential of the best-fit NFW profile instead of the numerical one. This more approximate method is the most efficient and may be of convenient use in analytical studies of haloes, for which the virial mass $M_{\mathrm{vir}}$ and concentration $c$ can be related to the parameters $\rho_{\mathrm{s}}$ and $r_{\mathrm{s}}$ of the NFW profile in a straightforward manner. While the NFW scaling relations have deficiencies at small energy due to significant deviations of the NFW potential from the simulated one (see Figs. 2 and 3), these deficiencies may not be relevant in many applications that concern halo regions at $r>0.05R_{\mathrm{vir}}$. Further, the central halo region is affected by baryonic processes, which cannot be addressed by models based on dark matter only.
Our scaling relations are inspired by those for the SIS and can account for the scaling law for the volumetric density of DF found by Arad et al. (2004). We have qualitatively discussed how the scaling relation for $dM/dE$ may arise during self-similar halo growth. We also note that the scaling of $f(E)$ with $r_{\mathrm{max}}(E)$ resembles the scaling law for the pseudo-phase-space density, $\rho(r)/\sigma^{3}(r)\propto r^{-\alpha}$ (Taylor & Navarro, 2001). The true origin of our scaling relations may shed important light on halo formation and merits further study. It is also interesting to go beyond the isotropic halo models and investigate if anisotropic ones follow any scaling relations.
@@ Acknowledgments
This work was supported in part by the US Department of Energy under grant DE-FG02-87ER40328 and by a Grant-in-Aid from the University of Minnesota. ZL thanks Jun Zhang for discussion regarding the scaling law of Arad et al. (2004). ZL acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101109759 (“CuspCore”) and from the Israel Science Foundation Grant ISF 861/20. The figures were created with the colorblind friendly scheme developed by Petroff (2021).

### v2

@@ 6 Summary and Conclusions
Based on the comparisons in Figs. 24 and Table 1, we suggest the following two efficient methods to approximate the energy distributions and DFs of haloes. For simulated haloes, one can use the scaling relations in Eqs. (20) and (21) after obtaining the best-fit NFW profile and the numerical potential. In more approximate contexts, one can use the potential of the best-fit NFW profile instead of the numerical one. This more approximate method is the most efficient and may be of convenient use in analytical studies of haloes, for which the virial mass $M_{\mathrm{vir}}$ and concentration $c$ can be related to the parameters $\rho_{\mathrm{s}}$ and $r_{\mathrm{s}}$ of the NFW profile in a straightforward manner. While the NFW scaling relations have deficiencies at small energy due to significant deviations of the NFW potential from the simulated one (see Figs. 2 and 3), these deficiencies may not be relevant in many applications that concern halo regions at $r>0.05R_{\mathrm{vir}}$. Further, the central halo region is affected by baryonic processes, which cannot be addressed by models which only include dark matter.
Our scaling relations are inspired by those for the SIS and can account for the scaling law for the volumetric density of DF found by Arad et al. (2004). We have qualitatively discussed how the scaling relation for $dM/dE$ may arise during self-similar halo growth. We also note that the scaling of $f(E)$ with $r_{\mathrm{max}}(E)$ resembles the scaling law for the pseudo-phase-space density, $\rho(r)/\sigma^{3}(r)\propto r^{-\gamma}$ (Taylor & Navarro, 2001). The true origin of our scaling relations may shed important light on halo formation and merits further study.
In general, dark matter haloes are only approximately spherical and have velocity anisotropy (see e.g., He et al. 2024). As mentioned in the introduction, the energy distribution $dM/dE$ of a spherical halo is mostly determined by its density profile and is insensitive to its velocity anisotropy (Binney & Tremaine, 2008; Baes & Dejonghe, 2021). Therefore, so long as spherical density profiles and the corresponding potentials provide good description of the simulated haloes, the scaling relation presented here for $dM/dE$ should also provide good description of these haloes. On the other hand, the DF of the simulated haloes has significant dependence on velocity anisotropy (e.g., Wojtak et al., 2008). Because the scaling relation presented here for the DF assumes no velocity anisotropy, this result can be applied only when approximation of isotropic haloes is sufficient. While it is challenging to model the energy distribution and DF of anisotropic haloes (but see e.g., Wojtak et al., 2008), it is worthwhile to investigate how the scaling relations are modified by velocity anisotropy. Such studies may provide insights into the structure and formation of anisotropic haloes, in addition to prescribing good approximations to properties of simulated haloes.
@@ Acknowledgments
This work was supported in part by the US Department of Energy under grant DE-FG02-87ER40328 and by a Grant-in-Aid from the University of Minnesota. ZL thanks Jun Zhang for discussion regarding the scaling law of Arad et al. (2004). ZL acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101109759 (“CuspCore”) and from the Israel Science Foundation Grant (ISF 861/20, 3061/21). The figures were created with the colorblind friendly scheme developed by Petroff (2021).
@@
Appendix A Fits for individual haloes
We estimate the halo-to-halo scatter of the scaling relations by fitting Eqs. (17) and (18) to individual haloes using the same procedure described in §3.2. In Fig. 5, we present the best-fit sets of $(\alpha,m)$ and $(\beta,n)$ for each of the 79 haloes in our sample as black crosses, along with the $1\sigma$ (68%) fitting uncertainties as blue ellipses. The mean best-fit parameters of the sample (orange pluses) are $(\alpha,m)=(0.603,1.01)$ and $(\beta,n)=(0.036,-2.08)$, nearly identical to those values (red pluses) obtained from fitting the median halo results in §3.2. The rms fitting uncertainties for individual haloes are $\delta=0.015$, 0.021, 0.0012, and 0.020 for $\alpha$, $m$, $\beta$, and $n$, respectively. In comparison, the standard deviations of the best-fit parameters are $\sigma=0.017$, 0.036, 0.0015, and 0.032, respectively. We quantify the intrinsic halo-to-halo scatter as $\sqrt{\sigma^{2}-\delta^{2}}=0.0084$, 0.029, 0.00090, and 0.024, respectively. These intrinsic scatters are only at the 1–3% level relative to the mean best-fit parameters. We thus conclude that the scaling relations obtained here are nearly universal across a wide range of haloes.
Figure 5: Left panel: Best-fit parameters $(\alpha,m)$ of Eq.  (17) for $dM/dE$ for each of the 79 haloes. Right panel: Best-fit parameters $(\beta,n)$ of Eq.  (18) for $f(E)$. In both panels, blue ellipses represent the $1\sigma$ (68%) fitting uncertainties around the best-fit parameters (black crosses) for individual haloes. Orange ellipses indicate intrinsic halo-to-halo scatters around the mean best-fit parameters (orange pluses), which nearly coincide with the red pluses for $(\alpha,m)=(0.60,1)$ and $(\beta,n)=(0.036,-2.08)$ obtained from fitting the median halo results in §3.2.